OpenStudy (anonymous):

g(x)=integral of f(t^2)

OpenStudy (turingtest):

do you know the fundamental theorem of calculus?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

i have up to now y=4(x-1)+g(1) I am not able to find g(1)

OpenStudy (anonymous):

First find \(g(x)\).

OpenStudy (anonymous):

g(x)=integral of f(t^2)2tdt from 0 to x=F(x^2)-F(0). No idea how to find the function F.

OpenStudy (anonymous):

You are not quite correct there.

OpenStudy (anonymous):

Or rather, you are not going about it the right way.

OpenStudy (anonymous):

I don't really know how that is wrong!

OpenStudy (anonymous):

We don't want to introduce \(F(x)\) in the equation since \(F(x) = g(x) + c\) which introduces more uncertainty into this problem.

OpenStudy (anonymous):

\[ g(x) = \int_0^xf(t^2)2t\;dt \]First consider \[ h(t) = f(t^2)2t \]Giving us: \[ g(x) = \int_0^xh(t)dt \]

OpenStudy (anonymous):

We know now \[ g'(x) = h(x) = f(x^2)2x \]

OpenStudy (anonymous):

this brings me back to H(x)-H(0).

OpenStudy (anonymous):

f(x)-f(0).

OpenStudy (anonymous):

Thus \[ g'(1) = f(1)\cdot 2(1) \]

OpenStudy (anonymous):

f(1)-f(0)=4-2

OpenStudy (anonymous):

g(1)=2

OpenStudy (anonymous):

yes, the slope of the tangent line is 4 But g(1)=? I think my previous steps are wrong.

OpenStudy (anonymous):

Anyone? Please help. I am lost.

OpenStudy (turingtest):

are you sure that g'(x) is f(x)2x wio? I think it's just f(x)

OpenStudy (anonymous):

There is one thing I'm trying to work out.

OpenStudy (anonymous):

@TuringTest of that, I'm sure. Even if you do that \(u=t^2\) sub, you will have to change your bounds to be from \(0\) to \(x^2\).

OpenStudy (turingtest):

oh yeah, sorry

OpenStudy (anonymous):

Finding \(g(1)\) is a bit tricky.

OpenStudy (anonymous):

Have you found a way though :O?

OpenStudy (anonymous):

I spent 3h on this.

OpenStudy (anonymous):

Hmmm \[ g(x) - g(1) = \int_1^x f(t^2)2t\:dt \]

OpenStudy (turingtest):

yep, that;s what I'm staring at lol

OpenStudy (anonymous):

I've also looked at that for a long time! No idea how to solve. I'm losing all hopes.

OpenStudy (anonymous):



OpenStudy (anonymous):

Well, since the problem assumes that \(f(t)\) is any function, why not try some sample functions?

OpenStudy (anonymous):

Try \(f(t) = 1\) and \(f(t) = 1\) and see what you come up with.

OpenStudy (anonymous):

I mean \(f(t) = 1\).

OpenStudy (anonymous):

Crap, I mean \(f(t)=t\).

OpenStudy (anonymous):

Actually, scratch that, all we need is to come up with a sample function that meets our conditions. Those conditions being: \(f(0)=4\) and \(f(1) = 2\).

OpenStudy (anonymous):

So try \(f(t) = 4- 2t\)

OpenStudy (anonymous):

you mean f(t^2)=4-2t^2?